OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4.

OpenStudy (anonymous):

Rearrange the line equation in term of : y = mx + c

OpenStudy (anonymous):

Firstly add 3y both the sides..

OpenStudy (anonymous):

Can you do this??

OpenStudy (anonymous):

idk how to do this im taking the pretest now..

OpenStudy (anonymous):

Then you need to study this first that is how you can learn something..

OpenStudy (anonymous):

yea my teachere is gonna teach the lesson tomorrow

OpenStudy (anonymous):

See when you will add 3y both the sides then you will get like this: \[3x - 3y = 4 \implies 3x - 3y + 3y = 4 + 3y\] Now cancel whatever is being cancelled: \[3x - \cancel{3y} + \cancel{3y} = 4 + 3y \implies 3x = 4 + 3y\]

OpenStudy (anonymous):

Now, subtract 4 both the sides: \[3x - 4 = 4 + 3y - 4 \implies 3x - 4 = \cancel{4} + 3y - \cancel{4} \implies 3y = 3x - 4\] Now divide it by 3: \[\frac{3y}{3} = \frac{3x}{3} -\frac{4}{3} \implies y = x - \frac{4}{3}\]

OpenStudy (anonymous):

i see alhough my options are confusing me

OpenStudy (anonymous):

slope = -1; y intercept at (0, -) slope = 1; y intercept at (0, -) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (anonymous):

Your question must be : 3x - 3y = 3 and not 4.. Check it..

OpenStudy (anonymous):

no sorry it says 4

OpenStudy (anonymous):

Then there must be something wrong.. tell me your b choice once again..

OpenStudy (anonymous):

i apoligise !

OpenStudy (anonymous):

slope = -1; y intercept at (0, -4/3) slope = 1; y intercept at (0, -4/3) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (anonymous):

so sorry i didnt notice

OpenStudy (anonymous):

Hey, no need to be..

OpenStudy (anonymous):

See we have got: y = x - 4/3 I can write it as: \[ y = x + (-\frac{4}{3})\] \[y = mx + c\] can you compare and tell me what is m and c here??

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

i dont think i can

OpenStudy (anonymous):

this is confusing i havent learned it all yet but i still dont want to get the answer wrong

OpenStudy (anonymous):

What you have to just compare them: see for colored part for easier comparison: \[y = \color{red}{m}x + \color{green}{c}\] \[y = \color{red}{1}{x} + \color{green}{\frac{4}{3}}\] Now tell me what are the red colored parts??

OpenStudy (anonymous):

oh ok thst helpd thank you

OpenStudy (anonymous):

Hey what happened..??

OpenStudy (anonymous):

so the answer would be b ?

OpenStudy (anonymous):

m = 1 and c = -4/3 yep answer is b..

OpenStudy (anonymous):

ah thank s!

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

can you help with one more

OpenStudy (anonymous):

i feel like im wasting time trying to figure it out when my teacher is gonna go step by step tomorrow

OpenStudy (anonymous):

It is already 3:53am here, so sorry, but I am sleepy now... Good Night Miami33..

OpenStudy (anonymous):

oh ok thanks again

OpenStudy (anonymous):

Nothing to mention... Keep Studying..